OpenStudy (anonymous):

how would I solve (x+2)^2=1??? And what would the answer be??

OpenStudy (anonymous):

Expand, then bring the 1 over the other side. Factor the this final quadratic and get the values of x.

OpenStudy (anonymous):

\[x ^{2}+4x+4=1\]

OpenStudy (anonymous):

\[x ^{2}+4x+3=0\]

OpenStudy (anonymous):

I got x=-1 and -3?

OpenStudy (anonymous):

\[(x+1)(x+3)=0\]

OpenStudy (anonymous):

cool @laurenh22

OpenStudy (anonymous):

now give me some love @laurenh22